Circular No. 6888

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998bn IN NGC 4462 F. Patat, European Southern Observatory (ESO); and M. Maia, Obser-

vatorio Nacional, Rio de Janeiro, write: “We observed SN 1998bn with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph; 5-min exposure, range 360–750 nm, resolution 0.5 nm) on Apr. 28.10 UT. A full reduction of the CCD spectrogram shows that this is a type-Ia supernova, ∼ 4 days before maximum light, very similar to the well-studied SN 1994D (Patat et al. 1996, MNRAS 278, 111). The spectrum is dominated by a multitude of lines with the usual P-Cyg profiles, superimposed on a blue continuum. Most of the lines arise from Si II, S II, Fe II, and Fe III. The expansion velocities deduced from the minima of Si II 635.5-nm and Ca II H and K are 11 500 and 15 500 km/s, respectively. An unresolved interstellar absorption of Na I D, arising in the parent galaxy, is clearly visible at 592.4 nm, with an equivalent width of 0.08 nm. The recession velocity deduced from this feature (1580 km/s) is slightly smaller than the value reported by Tully (1988, Nearby Galaxies Catalog) for NGC 4462 (1866 km/s).”

MARKARIAN 421 T. Takahashi, K. Maegami, F. Makino, and H. Ozawa, Institute of

Space and Astronautical Science; and H. Kubo, Institute of Physical and Chemical Research (RIKEN), on behalf of the international Markarian 421 campaign team, communicate: “During the intensive multifrequency campaign of this BL Lac object, ASCA detected a flare from the source on Apr. 27. The preliminary energy flux was 3.4 × 10^{−10} erg cm^{−2} s^{−1} (2– 10 keV), which is about the same as the peak flux that we observed in the 1994 x-ray/TeV flare (Takahashi et al. 1996, Ap.J. 470, L89). The photon index was ∼ 2.5. A 7-day continuous observation with ASCA began on Apr. 23. The preliminary analysis for the data taken since Apr. 23 shows that the source is quite bright (2.1 and 2.5 × 10^{−10} erg cm^{−2} s^{−1} on Apr. 24 and 25, respectively, at 2–10 keV). The reports from the CAT, HEGRA, and Whipple teams (based on their preliminary data analyses) tell us that Mrk 421 was very active during this period also in TeV γ-rays, with a strongly variable flux with an average value ∼ 1 Crab or so. Observations at other wavelengths are encouraged. For more details of the campaign schedule, send e-mail to takahasi@astro.isas.ac.jp; for current activities of the optical monitoring during the campaign, see

http://gamma.bu.edu/webt/mrk421.html.”

1998 April 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green